# Mind Meets Robots: A Review of EEG-Based Brain-Robot Interaction Systems

### v1

@@ 4.3.2. Evaluation Techniques
Dimension Sub-dimension Papers (%) References
Pure BCI
34(39.1%)
(Memar and Esfahani, 2018; Dissanayake et al., 2022; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Iwane et al., 2019; Lu et al., 2022; Alimardani et al., 2022, 2020; Long et al., 2019; Yoon et al., 2021; Chenga et al., 2023; Chu et al., 2018; Kilmarx et al., 2018; Wang et al., 2018c; Cao and Liu, 2018; Jo et al., 2022; Liu and Jebelli, 2021; Roshdy et al., 2021; Sanguantrakul et al., 2020; Penaloza et al., 2018; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Zhao et al., 2020b; Guo et al., 2020; Karunasena et al., 2021; Jiang et al., 2018; Quiles et al., 2022; Prinsen et al., 2022; Braun et al., 2019; Li et al., 2023; Farmaki et al., 2022; Chhabra et al., 2020; Wu et al., 2022; Sugiyama et al., 2023)
BCI + Agents
17(19.5%)
(Toichoa Eyam et al., 2021; Si-Mohammed et al., 2018; Chen et al., 2020; Kar et al., 2021; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Aznan et al., 2019; Li et al., 2019; Rashid et al., 2020; Ak et al., 2022; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Wang et al., 2018b; Ai et al., 2018; Fang et al., 2022; Lyu et al., 2022)
Proactive Control Single Signal Control (SSC)
47(54%)
(Belkacem and Lakas, 2021; Ehrlich and Cheng, 2019; Memar and Esfahani, 2018; Dissanayake et al., 2022; Toichoa Eyam et al., 2021; Iwane et al., 2019; Memar and Esfahani, 2019; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Chu et al., 2018; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Li et al., 2022a; Roy and Bhaumik, 2022; Liu and Jebelli, 2021; Martínez-Cagigal et al., 2020; Lu et al., 2020; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Roshdy et al., 2021; Sanguantrakul et al., 2020; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Rashid et al., 2020; Ak et al., 2022; Zhao et al., 2020b; Guo et al., 2020; Karunasena et al., 2021; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Ai et al., 2018; Quiles et al., 2022; Li et al., 2023; Kim et al., 2021; Farmaki et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Wu et al., 2022; Sugiyama et al., 2023)
Hybrid Signal Control (HSC)
9(10.3%)
(Lu et al., 2022; Tang et al., 2022; Pawuś and Paszkiel, 2022; Wang et al., 2018c; Jo et al., 2022; Li et al., 2019; Gordleeva et al., 2020; Nann et al., 2020; Korovesis et al., 2019)
Task-oriented HRI
17(19.5%)
(Baka et al., 2019; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Kompatsiari et al., 2018; Alimardani et al., 2022, 2020; Aldini et al., 2022; Aznan et al., 2019; Penaloza et al., 2018; Yu et al., 2019; Prinsen et al., 2022; Braun et al., 2019; Qian et al., 2018; Richter et al., 2023)
@@ 5.3.2. Hybrid Signal Control
LITERATURE
BRAIN
Signal Acquisition
BRAIN
Signal Decoding
INTERACTION ROBOT
@@
LITERATURE
BRAIN
Signal Acquisition
BRAIN
Signal Decoding
INTERACTION ROBOT

### v3

@@ 5. Analysis of The Interaction Entity
Table 5. The evaluation methods identified in our corpus.
Metrics Papers (%) References
Evaluation Methods Performance metrics 54 (62.1%) (Belkacem and Lakas, 2021; Memar and Esfahani, 2018; Staffa and Rossi, 2022; Iwane et al., 2019; Kompatsiari et al., 2018; Long et al., 2019; Si-Mohammed et al., 2018; Chenga et al., 2023; Tang et al., 2022; Pawuś and Paszkiel, 2022; Kar et al., 2021; Chu et al., 2018; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Jo et al., 2022; Aldini et al., 2022; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Roshdy et al., 2021; Aznan et al., 2019; Sanguantrakul et al., 2020; Penaloza et al., 2018; Yu et al., 2019; Bahman and Shamsollahi, 2019; Rashid et al., 2020; Zhao et al., 2020b; Guo et al., 2020; Nann et al., 2020; Gordleeva et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Kuffuor and Samanta, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Quiles et al., 2022; Mondini et al., 2020; Wei et al., 2021; Araujo et al., 2021; Rahul and Sharma, 2019; Francis et al., 2021; Shao et al., 2020; Xu et al., 2018; Boonarchatong and Ketcham, 2023; Prinsen et al., 2022; Li et al., 2023; Qian et al., 2018; Farmaki et al., 2022; Lyu et al., 2022; Wu et al., 2022; Sugiyama et al., 2023)
UX metrics 16 (18.4%) (Baka et al., 2019; Memar and Esfahani, 2018; Chang and Sun, 2021; Toichoa Eyam et al., 2021; Staffa and Rossi, 2022; Alimardani et al., 2020; Chenga et al., 2023; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Martínez-Cagigal et al., 2020; Nann et al., 2020; Li et al., 2023; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Richter et al., 2023)
Signal information 20 (23%) (Ehrlich and Cheng, 2019; Dissanayake et al., 2022; Aldini et al., 2019; Wang and Sugaya, 2021; Lu et al., 2022; Alimardani et al., 2022; Chenga et al., 2023; Chu et al., 2018; Li et al., 2022a; Roy and Bhaumik, 2022; Aldini et al., 2022; Penaloza et al., 2018; Yu et al., 2019; Li et al., 2019; Chiuzbaian et al., 2019; Karunasena et al., 2021; Yuan and Li, 2018; Ai et al., 2018; Mondini et al., 2020; Kim et al., 2021)
@@ 5.2. BCI+Agents
Dimension Sub-dimension Papers (%) References
Pure BCI 34(39.1%) (Memar and Esfahani, 2018; Dissanayake et al., 2022; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Iwane et al., 2019; Lu et al., 2022; Alimardani et al., 2022, 2020; Long et al., 2019; Yoon et al., 2021; Chenga et al., 2023; Chu et al., 2018; Kilmarx et al., 2018; Wang et al., 2018c; Cao and Liu, 2018; Jo et al., 2022; Liu and Jebelli, 2021; Roshdy et al., 2021; Sanguantrakul et al., 2020; Penaloza et al., 2018; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Zhao et al., 2020b; Guo et al., 2020; Karunasena et al., 2021; Jiang et al., 2018; Quiles et al., 2022; Prinsen et al., 2022; Braun et al., 2019; Li et al., 2023; Farmaki et al., 2022; Chhabra et al., 2020; Wu et al., 2022; Sugiyama et al., 2023)
BCI + Agents 17(19.5%) (Toichoa Eyam et al., 2021; Si-Mohammed et al., 2018; Chen et al., 2020; Kar et al., 2021; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Aznan et al., 2019; Li et al., 2019; Rashid et al., 2020; Ak et al., 2022; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Wang et al., 2018b; Ai et al., 2018; Fang et al., 2022; Lyu et al., 2022)
Proactive Control Single Signal Control (SSC) 47(54%) (Belkacem and Lakas, 2021; Ehrlich and Cheng, 2019; Memar and Esfahani, 2018; Dissanayake et al., 2022; Toichoa Eyam et al., 2021; Iwane et al., 2019; Memar and Esfahani, 2019; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Chu et al., 2018; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Li et al., 2022a; Roy and Bhaumik, 2022; Liu and Jebelli, 2021; Martínez-Cagigal et al., 2020; Lu et al., 2020; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Roshdy et al., 2021; Sanguantrakul et al., 2020; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Rashid et al., 2020; Ak et al., 2022; Zhao et al., 2020b; Guo et al., 2020; Karunasena et al., 2021; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Ai et al., 2018; Quiles et al., 2022; Li et al., 2023; Kim et al., 2021; Farmaki et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Wu et al., 2022; Sugiyama et al., 2023)
Hybrid Signal Control (HSC) 9(10.3%) (Lu et al., 2022; Tang et al., 2022; Pawuś and Paszkiel, 2022; Wang et al., 2018c; Jo et al., 2022; Li et al., 2019; Gordleeva et al., 2020; Nann et al., 2020; Korovesis et al., 2019)
Task-oriented HRI 17(19.5%) (Baka et al., 2019; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Kompatsiari et al., 2018; Alimardani et al., 2022, 2020; Aldini et al., 2022; Aznan et al., 2019; Penaloza et al., 2018; Yu et al., 2019; Prinsen et al., 2022; Braun et al., 2019; Qian et al., 2018; Richter et al., 2023)
@@ 5.3.2. Hybrid Signal Control
LITERATURE BRAIN Signal Acquisition BRAIN Signal Decoding INTERACTION ROBOT
@@
LITERATURE BRAIN Signal Acquisition BRAIN Signal Decoding INTERACTION ROBOT